# CerberusDet: Unified Multi-Task Object Detection

### v1

@@ Abstract
Object detection is a core task in computer vision. Over the years, the development of numerous models has significantly enhanced performance. However, these conventional models are usually limited by the data on which they were trained and by the category logic they define. With the recent rise of Language-Visual Models, new methods have emerged that are not restricted to these fixed categories. Despite their flexibility, such Open Vocabulary detection models still fall short in accuracy compared to traditional models with fixed classes. At the same time, more accurate data-specific models face challenges when there is a need to extend classes or merge different datasets for training. The latter often cannot be combined due to different logics or conflicting class definitions, making it difficult to improve a model without compromising its performance. In this paper, we introduce CerberusDet, a framework with a multi-headed model designed for handling multiple object detection tasks. Proposed model is built on the YOLO architecture and efficiently shares visual features from both backbone and neck components, while maintaining separate task heads. This approach allows CerberusDet to perform very efficiently while still delivering optimal results. We evaluated the model on the PASCAL VOC dataset and additional categories from the Objects365 dataset to demonstrate its abilities. CerberusDet achieved results comparable to state-of-the-art data-specific models with 36% less inference time. The more tasks are trained together, the more efficient the proposed model becomes compared to running individual models sequentially. The training and inference code, as well as the model, are available as open-source. 111https://github.com/ai-forever/CerberusDet
@@ 1 Introduction
The goal of this work is to build a unified model trained on multiple datasets that does not degrade in accuracy compared to individually trained models, while utilizing less computational budget. We present the CerberusDet framework for training a single detection neural network on multiple datasets simultaneously. We also demonstrate an approach to identify the optimal model architecture, as not all tasks can be trained together. A notable challenge lies in determining which parameters to share across which tasks. Suboptimal grouping of tasks may cause negative transfer [37], the problem of sharing information between unrelated tasks. Additionally, with computational resource constraints, proposed approach allows us to choose the architecture that fits the requirements. To evaluate the proposed CerberusDet model, we conduct experiments with open data and obtain results comparable to separated data-specific state-of-the-art models, but with one unified neural network. The presented architecture is based on YOLO[40], but the method can be easily adapted to many other OD architectures.”
@@
We present experimental results using open datasets, offering insights into the effectiveness of various approaches.
@@ 3.1 Method
The CerberusDet model is built upon the YOLO [40] architecture. It optimizes computational resources by sharing all backbone parameters across tasks, while each task retains its own unique set of parameters for the head. The neck layers can either be shared or specific to the task. An example of YOLOv8-based architecture for three tasks is illustrated in Figure 2. With the standard YOLOv8x architecture and 640 input image resolution, the model backbone consists of 184 layers and 30M parameters. The neck has 6 shareable modules with 134 layers and 28M parameters. Each head consists of 54 layers and 8M parameters.
@@ 3.4 The impact of training settings
All models were built on top of YOLOv5x with an input image resolution of 640x640, measurements were made with fp16 precision on the V100 GPU.
@@ 4 Open-source datasets experiments
Within this section, we outline CerberusDet’s experimental setup, results, and training configuration. Furthermore, we conduct a comparative analysis between CerberusDet and standalone YOLOv8 models using public datasets. This comparison also involves the previous best-performing model, Cascade Eff-B7 NAS-FPN [18] trained with the PASCAL VOC dataset.
Our analysis includes the presentation of metrics such as mean average precision (mAP): mAP@0.5 and mAP@0.5:0.95, alongside measurements of the models inference speeds.
@@ 4.1 Datasets
We conducted experiments on two publicly available datasets: PASCAL VOC and Objects365. Objects365 was filtered to include only images containing specific animal objects not present in the PASCAL VOC dataset. This resulted in two datasets, both balanced in terms of class distribution and image count.
@@
In both cases, standard datasets with predefined splits are used. For Objects365, these include the train and val sets, while for PASCAL VOC, they encompass train/val/test for the year 2007 and train/val for the year 2012. To filter Objects365, all non-empty images with annotations containing at least one object belonging to the animal class were used. Empty images without annotations and objects from other classes were filtered out. A detailed list of classes and the filtering code are available in the project repository.
@@
The PASCAL VOC dataset comprises 20 classes with 16,551 images in the training set and 4,952 images in the validation set. For training, we combined the trainval sets of PASCAL VOC 2007 and PASCAL VOC 2012 and validated on the test set of PASCAL VOC 2007. The Objects365 dataset consists of 19 classes with 14,295 images in the training set and 5,413 images in the validation set.
| Model | Train Set | Test Set | mAP | Speed (ms) | |
| 0.5 IoU | 0.5:0.95 IoU | ||||
| Cascade Eff-B7 NAS-FPN [18] | VOC 2007 + VOC 2012 | VOC 2007 | 0.8930 | - | - |
| YOLOv8 | VOC 2007 + VOC 2012 | VOC 2007 | 5.6 | ||
| YOLOv8 | Objects365 animals | Objects365 animals | 5.6 | ||
| CerberusDet | VOC 2007 + VOC 2012 Objects365 animals | VOC 2007 | 7.2 | ||
| Objects365 animals | |||||
4.2 Model Training
@@
We trained CerberusDet on the aforementioned datasets concurrently using hyperparameters inherited from YOLOv8, which were fine-tuned via hyperparameter search. CerberusDet leverages the YOLOv8x as its backbone architecture. The vanilla YOLOv8 head is used for each detection task, all neck parameters are task-specific. Additionally, we utilized transfer learning by initializing CerberusDet’s backbone and neck with a pre-trained model on COCO by the authors of YOLOv8, while the heads parameters were initialized randomly.
@@
The training was completed in 27 hours using 1 Nvidia V100 GPU with a batch size of 32. The training was done with mixed precision, SGD optimizer and synchronized batch normalization. The input resolution was set to 640x640 pixels.
@@ 4.3 Experimental Results
Experiments yielded promising results in terms of mAP scores (see Table 2), comparable to state-of-the-art separated data-specific model. CerberusDet demonstrated superior performance, achieving higher mAP@0.5 scores compared to standalone YOLOv8 models for both datasets.
Importantly, CerberusDet confidently surpasses the previous best-performing model on the PASCAL VOC dataset by a significant margin in terms of the mAP0.5 metric.
Additionally, even though CerberusDet performs inference for two tasks, it remains efficient, completing inference in 7.2 milliseconds on a single NVIDIA V100 GPU with float16 numerical precision. This is faster than inference using separate YOLOv8 models, which take 11.2 milliseconds – equivalent to 5.6 milliseconds per model.
The results demonstrate effectiveness across diverse datasets, surpassing standalone YOLOv8 models in terms of mAP@0.5 metric and computational efficiency.
@@ 5 Conclusions
Furthermore, CerberusDet is designed to be easily expandable, supporting additional tasks beyond object detection, such as attribute recognition and embedding calculation. While experiments utilized the YOLOv8 model, the method can be adapted to any object detection model architecture.
@@
6 Limitations
When training a model on multiple tasks, it is crucial to identify which tasks can share more parameters and which require fewer shared parameters. To address this, we provide implementation of the RSA algorithm, which, while requiring models trained for individual tasks, is still faster than iteratively testing all possible architecture variations with different parameter-sharing schemes. However, if training such models is not feasible, our approach can still be applied using a parameter-sharing scheme where all neck layers are task-dependent. While this may not be the most optimal solution, it remains more efficient than sequentially inferring single models.

### v2

# CerberusDet: Unified Multi-Dataset Object Detection

@@ Abstract
Conventional object detection models are usually limited by the data on which they were trained and by the category logic they define. With the recent rise of Language-Visual Models, new methods have emerged that are not restricted to these fixed categories. Despite their flexibility, such Open Vocabulary detection models still fall short in accuracy compared to traditional models with fixed classes. At the same time, more accurate data-specific models face challenges when there is a need to extend classes or merge different datasets for training. The latter often cannot be combined due to different logics or conflicting class definitions, making it difficult to improve a model without compromising its performance. In this paper, we introduce CerberusDet, a framework with a multi-headed model designed for handling multiple object detection tasks. Proposed model is built on the YOLO architecture and efficiently shares visual features from both backbone and neck components, while maintaining separate task heads. This approach allows CerberusDet to perform very efficiently while still delivering optimal results. We evaluated the model on the PASCAL VOC dataset and Objects365 dataset to demonstrate its abilities. CerberusDet achieved state-of-the-art results with 36% less inference time. The more tasks are trained together, the more efficient the proposed model becomes compared to running individual models sequentially. The training and inference code, as well as the model, are available as open-source. 111https://github.com/ai-forever/CerberusDet
@@ 1 Introduction
The goal of this work is to build a unified model trained on multiple datasets that does not degrade in accuracy compared to individually trained models, while utilizing less computational budget. We present the CerberusDet framework for training a single detection neural network on multiple datasets simultaneously. We also demonstrate an approach to identify the optimal model architecture, as not all tasks can be trained together. A notable challenge lies in determining which parameters to share across which tasks. Suboptimal grouping of tasks may cause negative transfer [41], the problem of sharing information between unrelated tasks. Additionally, with computational resource constraints, proposed approach allows us to choose the architecture that fits the requirements. To evaluate the proposed CerberusDet model, we conduct experiments with open data and obtain results comparable to separated data-specific state-of-the-art models, but with one unified neural network. The presented architecture is based on YOLO[44].”
@@
We present several experimental results using open datasets, offering insights into the effectiveness of various approaches.
@@ 3.1 Method
The CerberusDet model is built upon the YOLO [44] architecture. It optimizes computational resources by sharing all backbone parameters across tasks, while each task retains its own unique set of parameters for the head. The neck layers can either be shared or specific to the task. One of the possible variants of YOLOv8-based CerberusDet architecture for three tasks is illustrated in Figure 2. With the standard YOLOv8x architecture and 640 input image resolution, the model backbone consists of 184 layers and 30M parameters. The neck has 6 shareable modules with 134 layers and 28M parameters. Each head consists of 54 layers and 8M parameters.
@@ 3.4 The impact of training settings
Mentioned models were built on top of YOLOv5x with an input image resolution of 640x640, measurements were made with FP16 precision on the V100 GPU.
@@ 4 Open-source datasets experiments
Within this section, we outline CerberusDet’s experimental setup, results, and training configuration. Additionally, we perform a comparative analysis that includes CerberusDet and standalone YOLOv8 models using public datasets.
Notably, our comparison also incorporates the Open Vocabulary Detector YOLO-Worldv2-X [9], which was trained on a comprehensive set of base classes from the Objects365, GoldG [21], and CC3M [40] datasets at a resolution of 640. This inclusion enables us to evaluate the performance of a model employing a zero-shot approach for class expansion and comparison with traditional detection models.
Our analysis includes the presentation of metrics such as mean average precision (mAP): mAP@0.5 and mAP@0.5:0.95, alongside measurements of the models inference speed with FP16 precision on a V100 GPU.
@@ 4.1 Datasets
We conducted experiments on two publicly available datasets: PASCAL VOC [17] and Objects365 [39]. We conducted two experiments. The first was done with the original datasets, and the second used the PASCAL VOC dataset and two Objects365 subsets to simulate training on 3 datasets with different domains. The first subset was selected to include only images containing specific animal categories not present in the PASCAL VOC dataset. The second subset contains objects from various tableware categories.
@@
In both cases, standard datasets with predefined splits are used. For Objects365, these include the train and val sets, while for PASCAL VOC, they encompass train/val/test for the year 2007 and train/val for the year 2012. To create an Objects365 subset, all non-empty images with annotations containing at least one object belonging to the specific class were used. Empty images without annotations or images with objects from unrelated classes were filtered out. A detailed list of classes and the filtering code are available in the project repository.
@@
The PASCAL VOC dataset comprises 20 classes with 16,551 images in the training set and 4,952 images in the validation set. For training, we combined the trainval sets of PASCAL VOC 2007 and PASCAL VOC 2012, and validated on the test set of PASCAL VOC 2007. The original Objects365 dataset consists of objects from 365 classes, with 1.742M images in the training set and 79,578 images in the validation set. The Objects365 animals and Objects365 tableware subsets contain 19 and 12 classes, respectively. The training sets include 14,295 and 269,675 images respectively, while the validation sets include 5,413 and 24,645 images, respectively.
Model | Train Set | Test Set | mAP 0.5 | mAP 0.5:0.95 | Speed(ms) |
YOLOv8 | VOC 2007/VOC 2012 | VOC 2007 | 5.6 | ||
YOLOv8 | O365 | O365 | 5.6 | ||
YOLOv8 | O365 animals | O365 animals | 5.6 | ||
YOLOv8 | O365 tableware | O365 tableware | 5.6 | ||
YOLO-Worldv2-X† | O365 + GoldG + CC3M-Lite | VOC 2007 | 0.83 | 0.69 | 7.7 |
O365 | 0.31 | 0.24 | |||
O365 animals | 0.37 | 0.3 | |||
O365 tableware | 0.50 | 0.40 | |||
CerberusDet | VOC 2007/VOC 2012 + O365 | VOC 2007 | 7.2 | ||
O365 | |||||
CerberusDet | VOC 2007/VOC 2012 + O365 animals | VOC 2007 | 7.2 | ||
O365 animals | |||||
CerberusDet | VOC 2007/VOC 2012 + O365 animals + O365 tableware | VOC 2007 | 10 | ||
O365 animals | |||||
O365 tableware |
4.2 Models Training
Initially, we trained 4 individual YOLOv8x models on each of the datasets mentioned in section 4.1. These models were trained with a 640x640 input resolution, a batch size of 32, and mixed precision, serving as a baseline.
Then we trained the CerberusDet model on two datasets, Objects365 and PASCAL VOC, on datasets whose sizes differ by a factor of 100. We slightly modified the training procedure described in section 3.3 so that during each epoch, the model saw each dataset only once. To achieve this, we skipped batches from the smaller VOC dataset. Thus, in each iteration, the model saw samples from the Objects365 dataset, while samples from the VOC dataset were fed into the model only every 104 iteration. The batch size was set to 40. The model converged in 38 epochs.
We also trained the CerberusDet model on three datasets: Objects365 tableware, Objects365 animals, and PASCAL VOC. Similarly, there was a significant imbalance in the amount of data between the first and the other datasets. For training this model, we also modified the training process so that in each iteration, we used a different number of training examples from each dataset: 40, 4, and 4, respectively. The model was converged in 29 epochs.
We also trained the CerberusDet model on two datasets: PASCAL VOC and Objects365 animals to assess how the metrics change when a new dataset is added. We used standard training procedure, batch size of 32 and one V100 GPU to train this model.
@@ 4.2 Models Training
We used the same set of hyperparameters for all training procedures, inherited from YOLOv8x single models training. CerberusDet models leverage the YOLOv8x architecture as the backbone. The vanilla YOLOv8 head is used for each detection task, with all neck parameters being dataset-specific. Additionally, we utilized transfer learning by initializing CerberusDet’s backbone and neck with a pre-trained model on COCO povided by the YOLOv8 authors, while the heads parameters were initialized randomly.
@@
CerberusDet models were trained with mixed precision, the SGD optimizer and synchronized batch normalization using 8 H100 GPUs. The input resolution was set to 640x640 pixels.
@@ 4.3 Experimental Results
The experimental results are presented in Table 2.
We can observe that all three CerberusDet models either achieve the same accuracy as the individual YOLOv8 models or outperform them. When comparing the CerberusDet trained on two datasets (VOC, O365 animals) with the one trained on three (VOC, O365 animals, O365 tableware), we can see that adding another dataset to the training has minimal impact on the accuracy for the first two datasets.
In comparison with YOLO-Worldv2-X, a specialized CerberusDet models demonstrate a significant advantages. This comparison confirms the hypothesis that OVD detectors, while flexible and capable of rapidly adapting to new classes, generally fall short of the accuracy achieved by specialized detectors like CerberusDet.
When evaluating model inference speed, CerberusDet remains efficient, performing inference for two datasets in 7.2 milliseconds on a single NVIDIA V100 GPU with FP16 precision, faster than separate YOLOv8 models (11.2 ms). Extending CerberusDet to three datasets increases inference time to 10 milliseconds, which is still one-third faster than using three separate models. For two tasks, CerberusDet also shows a slight advantage over YOLO-Worldv2-X, with 7.2 ms compared to 7.7 ms.
The results demonstrate effectiveness of the CerberusDet across diverse datasets, even when a significant imbalance in the amount of data is present. Results also highlights an advantage over the zero-shot approach while maintaining flexibility in class expansion.
5 Limitations
When training a model on multiple tasks, it is important to identify which tasks can share more parameters and which require fewer shared parameters. To address this, we use the RSA algorithm, which, while requiring models trained for individual tasks, is still faster than iteratively testing all possible architecture variations with different parameter-sharing schemes. However, if training such models is not feasible, our approach can still be applied using a parameter-sharing scheme where all neck layers are task-dependent. While this may not be the most optimal solution, it remains more efficient than sequentially inferring single models.
@@ 6 Conclusions
Experiments with the zero-shot approach showed that using OVD in a similar scenario across multiple datasets () does not provide an advantage. In key metrics, especially those important in real-world applications such as speed and, most importantly, accuracy, OVD performs worse. This gap is explained by the need for a large amount of data for training, including target data, and the tendency to overfit on base classes. The only real advantage of the zero-shot approach in the current scenario is its speed and flexibility in adding potentially new and infinite categories, which can be useful in developing and testing class schemes. In practice, CerberusDet can be viewed as an intermediate solution between traditional detectors and OVD, taking the best of both.
Furthermore, CerberusDet is designed to be easily expandable, supporting additional tasks beyond object detection, such as attribute recognition and embedding calculation.